OpenStudy (anonymous):

How do you find the G(40) of this equation? G(x)=log sub3 (2x+1)-2 ?

OpenStudy (anonymous):

well this is how the problem was solved i'm just trying to figure out what happend to the log3 and the (81) towards the end. G(40)= log sub3(2(40)+1)-2 = log sub3 (81)-2 G(40)= 2

OpenStudy (anonymous):

\(\large G(40)=log_3(2\cdot 40+1)-2 \) \(\large G(40)=log_3(80+1)-2 \) \(\large G(40)=log_3(81)-2 \) now, \(\large log_3(81)=m \Leftrightarrow 3^m=81 \) this means m=4 so \(\large G(40)=4-2=2 \) .....:)

OpenStudy (anonymous):

thank you